ganeshie8 (ganeshie8):

Let \(R_n=111\ldots\text{n 1's}\) be the number with \(n\) ones in decimal number system. Find the smallest \(R_n\) that is divisible by \(83\).

OpenStudy (dan815):

.

OpenStudy (anonymous):

Correction: \[R_n=\sum_{k=0}^{\color{red}{n-1}}10^k=\frac{10^{\color{red}n}-1}{9}\] so find \(n\) such that \[\frac{10^{\color{red}n}-1}{9}\equiv0\mod83\]

OpenStudy (anonymous):

I suppose this can be simplified a bit... \(n\) should also satisfy this, no? \[10^n\equiv1\mod83\]

OpenStudy (dan815):

from fermats little theorem we have this bount atleast 10^(82)-1=0 mod 83

OpenStudy (dan815):

oh we can reduce the ones we have left to check by looking at the divisors of 82 apparently

OpenStudy (dan815):

acording to this http://prntscr.com/8834ha

OpenStudy (dan815):

82 = 2*41 therefore all we have left to check is 2,41 10^2-1 =/=0mod83 10^41-1 = 0 mod 83

OpenStudy (dan815):

41, 1s

OpenStudy (anonymous):

I'm getting the same result with Mathematica, checks out

OpenStudy (dan815):

the 2nd statement that there msut exist some lower exponent d that divides p-1, if p does not divide a looks interesting

OpenStudy (ikram002p):

i has same method finding the first n which satisfy 10^n=1 mod 83

OpenStudy (ikram002p):

so we know 10^82 satisfy, now we check its divisors 1,2,41,82 the end xD

ganeshie8 (ganeshie8):

Awesome! so it seems, given a prime, \(p\), we can always find the smallest \(R_n\) divisible by \(p\), by looking at the divisors of \(p-1\)

OpenStudy (jhannybean):

.

OpenStudy (ikram002p):

@ganeshie8 are you saying that R_n always composite ?

ganeshie8 (ganeshie8):

11 isn't composite

OpenStudy (ikram002p):

i know it has some primes im trying to understand what u have said :)

ganeshie8 (ganeshie8):

fix a prime \(p\) and form the set of all \(R_n\)'s that are divisible by \(p\). then we can find the least element in that set using the earlier trick

OpenStudy (ikram002p):

got it!

OpenStudy (ikram002p):

:O its awesome isnt it ?

ganeshie8 (ganeshie8):

so baiscally all \(\large R_{p-1}\)'s are composite because they are trivially divisible by \(p\)

OpenStudy (ikram002p):

yeah with even number of digits hmmm

ganeshie8 (ganeshie8):

Oh, all \(\large R_{2k}\) are composite is it ?

ganeshie8 (ganeshie8):

Easy to see that all \(\large R_{3k}\) are composite because they are divisible by \(3\)..

OpenStudy (ikram002p):

yeah for this one, i think we have proved this earlier

OpenStudy (ikram002p):

the 1010101 stuff

OpenStudy (ikram002p):

let me remember how it was like exactly

ganeshie8 (ganeshie8):

Ah, right, its easy one sec

ganeshie8 (ganeshie8):

\[\large R_{2k}=\sum_{k=0}^{\color{red}{2k-1}}10^k=\frac{10^{\color{red}{2k}}-1}{9} = \dfrac{(10^k-1)(10^k+1)}{9}\]

ganeshie8 (ganeshie8):

for \(k\gt 1\), we have \(10^k-1 = 9l\) such that \(l\gt 1\)

OpenStudy (ikram002p):

so ovs what l was whole number stays composite :3

ganeshie8 (ganeshie8):

for \(k\gt 1\), \[\large R_{2k}=\sum_{k=0}^{\color{red}{2k-1}}10^k=\frac{10^{\color{red}{2k}}-1}{9} = \dfrac{(10^k-1)(10^k+1)}{9} = l(10^k+1)\] which cannot be a prime as it is product of two factors that are > 1

OpenStudy (ikram002p):

haha sweet !

ganeshie8 (ganeshie8):

that also shows : if \(n\) is composite, then \(R_n\) is composite

OpenStudy (ikram002p):

correct!